# PoSh: Using Scene Graphs to Guide LLMs-as-a-Judge for Detailed Image Descriptions

### v2

@@ 6 Results & Discussion
On DOCENT, across all three dimensions, PoSh outperforms every existing replicable metric (i.e., metrics not reliant on an API), yielding a $0.11$ increase in Spearman $\rho$ for mistakes ($25\%\uparrow$), a $0.07$ increase for omissions ($14\%\uparrow$) and a $0.05$ increase for overall quality ($9\%\uparrow$) over the next best. It even outperforms GPT4o (in all settings) and text-only GPT5 (on omissions and overall quality). Among all metrics, DCScore (ye2025painting) proves best at predicting mistakes. However, its reliance on GPT4o to extract and verify factoids fails to achieve full coverage of reference detail, underperforming in predicting omissions and overall quality. Despite employing a smaller LLM, PoSh’s use of dependency parsing and coreference resolution to extract scene graphs avoids this.
On CapArena, PoSh achieves higher caption-level accuracies and model-ranking correlations than nearly every existing open-weight metric and GPT4o. The sole exception is LLaVa Critic, a much larger VLM-as-a-Judge (xiong2024llavacritic). This is driven in part by the simplicity of CapArena (see Table 1). On the subset of CapArena depicting three or more people (167 judgments), each of whom requires careful description, PoSh outperforms LLaVa Critic with model ranking correlations of $\mathbf{\rho=0.727,\tau=0.581}$ compared to $\mathbf{\rho=0.686,\tau=0.550}$. Thus, PoSh is robust to image type, excelling in visually complex cases that are of particular interest in detailed image description.
@@
In Table 7 of the Appendix, we report annotator agreement and aggregate preferences between SFT and DAPO with PoSh. In each dimension of interest, a PoSh-tuned generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. While PoSh-tuned generations had more mistakes (an average score of 0.243-0.243), these were incurred in service of much fewer missing details ($\mathbf{+0.432}$), resulting in higher overall quality ($\mathbf{+0.135}$). This speaks to the strength of PoSh when optimized directly. Given recent progress in generating synthetic detailed descriptions (li2024densefusion), PoSh-tuning could be freely scaled in post-training. Moreover, as PoSh produces localized granular scores, it supports token-level guidance (yang2023preference), an exciting direction to explore in future work.
@@
A core enabler of PoSh’s performance and interpretability is its thorough granular evaluation. It achieves this efficiently through inference optimizations like continuous batching and prefix caching (kwon2023efficient). PoSh scores the $400$ examples in DOCENT in $15$ minutes, or one every $2$ seconds, on a single H100 GPU. In contrast, DCScore (ye2025painting) takes upwards of $2$ hours due to its heavy use of GPT4 (one description every $25$ seconds). As manual evaluation takes $18$ minutes per description (see Table 1), PoSh effectively balances quality and cost.
@@ 9 Limitations
Recent efforts have explored using structural priors to guide generation (e.g. wang2025llava). As these methods extract and describe structure from images, and as PoSh extracts and validates structure from text, we do not expect PoSh to be biased towards them in an evaluation setting. Nevertheless, as these models become publicly available, this requires experimental validation.
@@ A.1.1 Comparison to Davidsonian Scene Graph
There are several differences between the Davidsonian Scene Graph (DSG) metric (chodavidsonian) and PoSh. While the emphasis of DSG is on evaluating text-to-image models, PoSh was designed specifically for detailed image descriptions and is tailored to the unique challenges they pose.
First, the emphasis of DSG is on evaluating text-to-image models: it compares an image to scene graph elements extracted from a prompt text through visual question answering. As such, it cannot serve as a reference-based metric for evaluating detailed image descriptions. Allowing the use of references is important. Downstream tasks have context-specific requirements that can only be specified with references. This is especially true in accessibility (deutsch2022limitations).
@@ A.2.1 Generations
    llava-v1.6-mistral-7b-hf on HuggingFace (liu2024improved)
@@
    Molmo-7B-D-0924 on HuggingFace (deitke2024molmo)
@@
We use the same prompt (included below). For LLaVA-1.5-7B and Molmo-D-7B, we use nucleus sampling holtzmancurious with $p=0.9$ and a temperature of $0.7$.
@@ A.2.2 Avoiding Double Penalties
In kasai-etal-2022-transparent, after identifying an error in precision, the authors correct the error before annotating recall. This avoids doubly penalizing a description for errors in specificity which would unfairly favor more generic descriptions (which are only penalized once, for recall). We instruct our annotators to do the same.
@@ A.3.3 Granular Baselines
We extract all of the $4$-grams from each sentence of a generation and its reference, embed them using Qwen/Qwen3-Embedding-8B (reimers-2019-sentence-bert; yang2025qwen3) and then calculate the maximum pairwise similarities between generation $4$-grams and reference $4$-grams. Generation text spans and reference text spans with maximum pairwise similarity scores lower than $0.7$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for 4GramEmbed in Table 2.
@@
We extract all of the components (objects, attribute-object pairs, and object-relation-object triples) from the scene graphs of a generation and its reference extracted for PoSh in Section 3, embed them using Qwen/Qwen3-Embedding-8B (reimers-2019-sentence-bert; yang2025qwen3) and then calculate the maximum pairwise similarities between the generation components and the reference components. Generation components and reference components with maximum pairwise similarity scores lower than $0.8$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for SGEmbed in Table 2.
@@ A.3.5 Reinforcement Learning
    DAPO (yu2025dapo) with full parameter updates, implemented with TRL (vonwerra2022trl), using a learning rate of 1e61e-6, 2020 warmup steps, $8$ generations per sample (with a temperature of $1.0$ and $top_{p}=0.7$), $\epsilon=0.28$, $\beta=0$, and an effective batch size of $64$ for a single epoch, choosing the final checkpoint

### v3

@@ 6 Results & Discussion
On DOCENT, across all three dimensions, PoSh outperforms every existing replicable metric (i.e., metrics not reliant on an API), yielding a $0.11$ increase in Spearman $\rho$ for mistakes ($25\%\uparrow$), a $0.07$ increase for omissions ($14\%\uparrow$) and a $0.05$ increase for overall quality ($9\%\uparrow$) over the next best. It even outperforms GPT4o (in all settings) and text-only GPT5 (on omissions and overall quality). Among all metrics, DCScore (Ye et al., 2025) proves best at predicting mistakes. However, its reliance on GPT4o to extract and verify factoids fails to achieve full coverage of reference detail, underperforming in predicting omissions and overall quality. Despite employing a smaller LLM, PoSh’s use of dependency parsing and coreference resolution to extract scene graphs avoids this.
On CapArena, PoSh achieves higher caption-level accuracies and model-ranking correlations than nearly every existing open-weight metric and GPT4o. The sole exception is LLaVa Critic, a much larger VLM-as-a-Judge (Xiong et al., 2025). This is driven in part by the simplicity of CapArena (see Table 1). On the subset of CapArena depicting three or more people (167 judgments), each of whom requires careful description, PoSh outperforms LLaVa Critic with model ranking correlations of $\mathbf{\rho=0.727,\tau=0.581}$ compared to $\mathbf{\rho=0.686,\tau=0.550}$. Thus, PoSh is robust to image type, excelling in visually complex cases that are of particular interest in detailed image description.
@@
In Table 7 of the Appendix, we report annotator agreement and aggregate preferences between SFT and DAPO with PoSh. In each dimension of interest, a PoSh-tuned generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. While PoSh-tuned generations had more mistakes (an average score of 0.243-0.243), these were incurred in service of much fewer missing details ($\mathbf{+0.432}$), resulting in higher overall quality ($\mathbf{+0.135}$). This speaks to the strength of PoSh when optimized directly. Given recent progress in generating synthetic detailed descriptions (Li et al., 2024), PoSh-tuning could be freely scaled in post-training. Moreover, as PoSh produces localized granular scores, it supports token-level guidance (Yang et al., 2023), an exciting direction to explore in future work.
@@
A core enabler of PoSh’s performance and interpretability is its thorough granular evaluation. It achieves this efficiently through inference optimizations like continuous batching and prefix caching (Kwon et al., 2023). PoSh scores the $400$ examples in DOCENT in $15$ minutes, or one every $2$ seconds, on a single H100 GPU. In contrast, DCScore (Ye et al., 2025) takes upwards of $2$ hours due to its heavy use of GPT4 (one description every $25$ seconds). As manual evaluation takes $18$ minutes per description (see Table 1), PoSh effectively balances quality and cost.
@@ 9 Limitations
Recent efforts have explored using structural priors to guide generation (e.g. Wang et al. (2025)). As these methods extract and describe structure from images, and as PoSh extracts and validates structure from text, we do not expect PoSh to be biased towards them in an evaluation setting. Nevertheless, as these models become publicly available, this requires experimental validation.
@@ A.1.1 Comparison to Davidsonian Scene Graph
There are several differences between the Davidsonian Scene Graph (DSG) metric (Cho et al., 2024) and PoSh. While the emphasis of DSG is on evaluating text-to-image models, PoSh was designed specifically for detailed image descriptions and is tailored to the unique challenges they pose.
First, the emphasis of DSG is on evaluating text-to-image models: it compares an image to scene graph elements extracted from a prompt text through visual question answering. As such, it cannot serve as a reference-based metric for evaluating detailed image descriptions. Allowing the use of references is important. Downstream tasks have context-specific requirements that can only be specified with references. This is especially true in accessibility (Deutsch et al., 2022).
@@ A.2.1 Generations
    llava-v1.6-mistral-7b-hf on HuggingFace (Liu et al., 2024)
@@
    Molmo-7B-D-0924 on HuggingFace (Deitke et al., 2024)
@@
We use the same prompt (included below). For LLaVA-1.5-7B and Molmo-D-7B, we use nucleus sampling Holtzman et al. (2019) with $p=0.9$ and a temperature of $0.7$.
@@ A.2.2 Avoiding Double Penalties
In Kasai et al. (2022), after identifying an error in precision, the authors correct the error before annotating recall. This avoids doubly penalizing a description for errors in specificity which would unfairly favor more generic descriptions (which are only penalized once, for recall). We instruct our annotators to do the same.
@@ A.3.3 Granular Baselines
We extract all of the $4$-grams from each sentence of a generation and its reference, embed them using Qwen/Qwen3-Embedding-8B (Reimers and Gurevych, 2019; Yang et al., 2025) and then calculate the maximum pairwise similarities between generation $4$-grams and reference $4$-grams. Generation text spans and reference text spans with maximum pairwise similarity scores lower than $0.7$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for 4GramEmbed in Table 2.
@@
We extract all of the components (objects, attribute-object pairs, and object-relation-object triples) from the scene graphs of a generation and its reference extracted for PoSh in Section 3, embed them using Qwen/Qwen3-Embedding-8B (Reimers and Gurevych, 2019; Yang et al., 2025) and then calculate the maximum pairwise similarities between the generation components and the reference components. Generation components and reference components with maximum pairwise similarity scores lower than $0.8$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for SGEmbed in Table 2.
@@ A.3.5 Reinforcement Learning
    DAPO (Yu et al., 2025) with full parameter updates, implemented with TRL (von Werra et al., 2020), using a learning rate of 1e61e-6, 2020 warmup steps, 88 generations per sample (with a temperature of $1.0$ and $top_{p}=0.7$), $\epsilon=0.28$, $\beta=0$, and an effective batch size of $64$ for a single epoch, choosing the final checkpoint